OpenStudy (anonymous):

FREE MEDALS WILL BE GIVEN OUT INJ ONE MINUTE TO ANY ONE WHO HELPS ME IN GEOMETRY

OpenStudy (anonymous):

:P

OpenStudy (secret-ninja):

Then go to the Geometry section.

OpenStudy (anonymous):

I DONT WANT TO BC THERES NOBODY GONNA HELP

OpenStudy (secret-ninja):

Okay, just post your question and I can TRY to help.

OpenStudy (anonymous):

GIVEN triangles abc and pqr,m<b=3x+4 and m<q=8x-6 A. 22 B. 11 C. 10 D. 25

OpenStudy (secret-ninja):

All you have to do to salve this is plug in the choices for "x". For example: m<b=3x+4 3 x 22 + 4 = 70. so that gives you m<b=70 Now you try m<q=8x-6 8 x 22 - 6 = 170 So that gives you m<q=170 They are not the same, so you need to try the next one. Can you go on from here?

OpenStudy (anonymous):

i did that to all and none of them match

OpenStudy (secret-ninja):

That's odd... I don't really know where to go one from there. Try just closing this question and reposting. Someone else might help. :)

OpenStudy (anonymous):

k